gracewikle:

If tan2x = 4 - 2tanx, then tanx = _____. a.)-1 ± sqrt 5 b.)-1 ± sqrt 3 c.)1 ± sqrt 3 d.)1 ± sqrt 5

justjm:

Is it \(tan^2x\)? Then that would make sense.

Nnesha:

no wonder why I couldn't find the solution. Yes it is squared.

Nnesha:

\[\rm \tan^2x=4-2~tanx\] tanx=sinx/cosx and tan^2x=sec^2x-1 we get \[\rm \sec^2x-1=4-2 \frac{sinx}{cosx}\] secx=1/cosx \[\rm \frac{ 1 }{ cox^2(x) }-1=4-2\frac{sinx}{cosx}\] \[\frac{ 1-\cos^2x }{ \cos^2x }=4-\frac{ 2sinx }{ cosx }\] 1-cos^2x=sin^2x

Nnesha:

stupid me don't know why doing all that work..

Nnesha:

\[\rm \tan^2x=4-2tanx\] \[\rm \tan^2x+2tanx-4=0\] let tanx= y \[\rm y^2+2y-4=0\] solve for y